# GFM-RAG: Graph Foundation Model for Retrieval Augmented Generation

### v1

@@ A.2 Domain-specific RAG Datasets
Appendix B Baselines
In experiments, we compare with several widely used retrieval methods under three categories: (1) single-step naive methods: BM25 (Robertson & Walker, 1994), Contriever (Izacard et al., 2022), GTR (Ni et al., 2022), ColBERTv2 (Santhanam et al., 2022), RAPTOR (Sarthi et al., 2024), Proposition (Chen et al., 2024); (2) graph-enhanced methods: LightRAG (Guo et al., 2024), HippoRAG (Gutiérrez et al., 2024); (3) multi-step methods: IRCoT (Trivedi et al., 2023). The detailed introduction of the baselines is as follows.
@@ Appendix B Baselines
  • BM25 (Robertson & Walker, 1994) is a classic information retrieval method based on the probabilistic model that ranks a set of documents based on the query terms frequency appearing in each document.
  • Contriever (Izacard et al., 2022) trains a dense retriever with contrastive learning on a large-scale corpus to retrieve relevant documents for a given query.
  • GTR (Ni et al., 2022) develops a scale-up T5-based dense retriever that could generalize across different datasets and domains.
  • ColBERTv2 (Santhanam et al., 2022) is a state-of-the-art dense retriever that couples an aggressive residual compression mechanism with a denoised supervision strategy to simultaneously improve the retrieval quality.
  • RAPTOR (Sarthi et al., 2024) is an LLM-augmented retriever that recursively embeds, clusters, and summarizes chunks of text, constructing a tree with differing levels of summarization to enable accurate retrieval.
  • Proposition (Chen et al., 2024) enhances the performance of dense retrievers by leveraging LLMs to generate a natural language proposition that captures the essential information of the document.
@@
  • LightRAG (Guo et al., 2024) is an innovative graph-enhanced RAG method that incorporates graph structures into text indexing and retrieval, enabling efficient retrieval of entities and their relationships. It employs a dual-level retrieval system to gather both low-level and high-level knowledge for LLM generation.
  • HippoRAG (Gutiérrez et al., 2024) is a state-of-the-art, training-free graph-enhanced retriever that uses the Personalized PageRank algorithm to assess entity relevance to a query and performs multi-hop retrieval on a document-based knowledge graph. It can be directly applied to various datasets.
@@
  • IRCoT (Trivedi et al., 2023) is a powerful multi-step retrieval pipeline that integrates the retrieval with the chain-of-thought (CoT) reasoning of LLMs. It guides the retrieval with CoT and in turn using retrieved documents to improve CoT. IRCoT can be compatible with arbitrary retrievers to conduct multi-step retrieval and reasoning.
Appendix C Implementations and Training Details
C.1 Training Data Construction
We extract 60,000 samples from the training set of HotpotQA, MuSiQu, and 2Wiki to construct KG-indexes and conduct large-scale training. Specifically, we merge the candidate passages as the document corpus. In the KG-index construction, we use the GPT-4o-mini (OpenAI, 2024a) with the OpenIE prompts described in HippoRAG (Gutiérrez et al., 2024) to extract the entities, relations, and triples from the document corpus. Then, we use the ColBERTv2 (Santhanam et al., 2022) to conduct the entity resolution by computing the similarity between entities as
@@ C.2 Model Settings
C.3 Training Settings
In the unsupervised KG completion pre-training, the GFM is trained on the mixture of 60 constructed KG-indexes for 30,000 steps. Then, we conduct the supervised document retrieval fine-tuning on the labeled question-document pairs for 5 epochs. The weight $\alpha$ between losses is set to 0.3. We use AdamW optimizer, learning rate of 5e-4 with batch sizes of both training stages set to 4. Each batch contains only one KG-index and training samples associated to it, where we randomly sample from different KG-indexes during training. The model is trained on 8 NVIDIA A100s (80G) with 14 hours pre-training and 5 hours fine-tuning. The detailed settings are summarized in Table 7.
@@ C.3 Training Settings
Retrieval Fine-tuning $\alpha$ 0.3
@@
Appendix D Additional Experiments
@@
D.1 Effectiveness of Different Sentence Embeddings
In this section, we study the effectiveness of different sentence embeddings in the GFM. We compare the all-mpnet-v2 (SBERT, 2021), bge-large-en (Xiao et al., 2023), gte-Qwen2-1.5B-instruct and gte-Qwen2-7B-instruct (Li et al., 2023b) as well as NV-Embed-v2 (Lee et al., 2024). We download the official pre-trained model from the Huggingface [^2] https://huggingface.co/. The details of the models are shown in Table 8. From the results, we can observe that the performance variance between different sentence embeddings is relatively small, where the all-mpnet-v2 achieves the best performance with respect to 3 metrics. This indicates that GFM-RAG is not sensitive to the choice of sentence embedding models. In experiments, we use the all-mpnet-v2 as the default sentence embedding model due to its efficiency. However, it has relative smaller context-size (512) which limits the length of input text. We leave the exploration of larger context-size sentence embedding models (e.g., NV-Embed-v2 with 32k context) for future work.
Table 9: Effectiveness of pre-training and fine-tuning in GFM-RAG.
Sentence Embedding Model HotpotQA MuSique 2Wiki
@@ D.1 Effectiveness of Different Sentence Embeddings
GFM-RAG $w/o$ Fine-tune 21.0 32.8 18.3 25.9 44.6 53.4
GFM-RAG $w/o$ Pre-train 77.8 86.5 48.3 58.3 88.3 92.5
D.2 Effectiveness of Different Training Strategies
In this section, we study the effectiveness of the two training tasks used in GFM-RAG. We compare the performance by only conducting the unsupervised KG completion pre-training (GFM-RAG $w/o$ Fine-tune) and supervised document retrieval fine-tuning (GFM-RAG $w/o$ Pre-train). The results are shown in Table 9. The results show that removing the supervised document retrieval fine-tuning significantly decreases the performance of GFM-RAG. This highlights the importance of supervised fine-tuning, as it enables the model to understand users’ queries and better capture the relevance between questions and knowledge for improved retrieval. Meanwhile, the unsupervised KG completion pre-training also plays a crucial role in enhancing the model’s reasoning ability by learning the knowledge graph structure and reasoning patterns from the large-scale triples in KG-indexes.
@@ D.2 Effectiveness of Different Training Strategies
D.3 Effectiveness of Loss Weights
In this section, we examine the effectiveness of the weights assigned to the BCE loss and ranking loss in training GFM-RAG. We compare performance by varying the weight $\alpha$ between the two losses: ${\mathcal{L}}=\alpha{\mathcal{L}}_{\text{BCE}}+(1-\alpha){\mathcal{L}}_{\text{RANK}}$, with results presented in Table 10. The findings indicate that using only either the BCE loss or ranking loss leads to suboptimal performance ($\alpha=0~{}\text{or}~{}1$). The best performance occurs when $\alpha$ is set to 0.3, which aligns with previous studies (Lin et al., 2024) suggesting that a smaller weight for BCE loss is preferable when positive samples are rare in the training data.
D.4 Model Transferability
@@
GFM-RAG (fine-tune) 82.7 75.9 60.8 49.5 77.5 86.6
In this section, we evaluate GFM-RAG’s transferability by fine-tuning on the training split of each domain-specific dataset. As shown in 11, GFM-RAG performs well in zero-shot generalization, with further improvements achieved through fine-tuning. This highlights its transferability when adapted to domain-specific datasets.

### v2

@@ Appendix A Query-dependent GNNs for Multi-hop Reasoning and Retrieval
Appendix B Datasets
B.1 Multi-hop QA Datasets
Three multi-hop QA datasets are used in our experiments: HotpotQA [72], MuSiQue [63], and 2WikiMultiHopQA (2Wiki) [20]. We provide a brief overview of these datasets below.
  • HotpotQA [72] is a multi-hop QA dataset that requires reasoning over multiple documents to answer questions. The dataset consists of 97k question-answer pairs, where each question is associated with up to 2 supporting and several distracting documents. The questions are designed to be answerable using multiple pieces of information from the supporting documents.
  • MuSiQue [63] is a challenging multi-hop QA dataset with 25k 2-4 hop questions. It requires coherent multi-step reasoning to answer questions that span multiple documents.
  • 2WikiMultiHopQA (2Wiki) [20] is a multi-hop QA dataset that requires reasoning over multiple Wikipedia articles to answer questions. The dataset consists of 192k questions, which are designed to be answerable using information from 2 or 4 articles.
In experiments, we adhere to the official data split to obtain the training samples and follow existing methods [64, 16] to use the same 1,000 samples from each validation set to avoid data leakage. We merge the candidate passages as the document corpus for KG-index construction. The statistics of the training and test data are presented in Table˜5 and Table˜6, respectively.
B.2 Domain-specific RAG Datasets
To test the generalizability of GFM-RAG, we evaluate it on seven domain-specific RAG datasets [10] including, (1) biomedical: PubMedQA [25]; (2) customer support: DelucionQA [54], TechQA [4], ExpertQA [39], EManual [44]; (3) general knowledge: MS Marco [45], HAGRID [27]. We provide a brief overview of these datasets below.
  • PubMedQA [25] is a collection of PubMed research abstracts with corresponding questions paired with 4 abstract chunks.
  • DelucionQA [54] is a domain-specific RAG dataset leveraging Jeep’s 2023 Gladiator model manual as the source of knowledge, where each question is associated with 4 context documents and only 1 relevant passage.
  • TechQA [4] is a collection of real-world user questions posted on IBMDeveloper and DeveloperWorks forums, along with 10 technical support documents relating to each question.
  • ExpertQA [39] is a collection of curated questions from domain experts in various fields of science, arts, and law. The dataset also contains expert-curated passages relevant to each question.
  • EManual [44] is a question-answering dataset comprising consumer electronic device manuals and realistic questions about them composed by human annotators, where each question is related with up to 3 context documents.
  • MS Marco [45] is an open-domain question-answering dataset sourced from Bing search engine user query logs. Each question is associated with 10 context passages retrieved via Bing web search.
  • HAGRID [27] is a multi-lingual information retrieval dataset with questions and passages from MIRACL [75].
In experiments, we use test sets constructed by RAGBench [10] and merge all the candidate passages as document corpus for KG-index construction. The statistics of the test dataset are detailed in Table˜6.
Appendix C Baselines
In experiments, we compare with several widely used retrieval methods under three categories: (1) single-step naive methods: BM25 [53], Contriever [22], GTR [46], ColBERTv2 [55], RAPTOR [56], Proposition [6]; (2) graph-enhanced methods: GraphRAG (MS) [9], LightRAG [15], HippoRAG [16]; (3) multi-step methods: Adaptive-RAG [23], FLARE [24], and IRCoT [64]. The detailed introduction of the baselines is as follows.
@@ Appendix C Baselines
  • BM25 [53] is a classic information retrieval method based on the probabilistic model that ranks a set of documents based on the query terms frequency appearing in each document.
  • Contriever [22] trains a dense retriever with contrastive learning on a large-scale corpus to retrieve relevant documents for a given query.
  • GTR [46] develops a scale-up T5-based dense retriever that could generalize across different datasets and domains.
  • ColBERTv2 [55] is a state-of-the-art dense retriever that couples an aggressive residual compression mechanism with a denoised supervision strategy to simultaneously improve the retrieval quality.
  • RAPTOR [56] is an LLM-augmented retriever that recursively embeds, clusters, and summarizes chunks of text, constructing a tree with differing levels of summarization to enable accurate retrieval.
  • Proposition [6] enhances the performance of dense retrievers by leveraging LLMs to generate a natural language proposition that captures the essential information of the document.
@@
  • GraphRAG (MS) [9] is a graph-enhanced retrieval method originally proposed by Microsoft. It builds a graph structure from the document corpus and use hierarchical community detection to cluster the documents into communities and generate a summary for each community. The summary together with the original documents are retrieved by the retriever for LLM generation.
  • LightRAG [15] is an innovative graph-enhanced RAG method that incorporates graph structures into text indexing and retrieval, enabling efficient retrieval of entities and their relationships. It employs a dual-level retrieval system to gather both low-level and high-level knowledge for LLM generation.
  • HippoRAG [16] is a state-of-the-art, training-free graph-enhanced retriever that uses the Personalized PageRank algorithm to assess entity relevance to a query and performs multi-hop retrieval on a document-based knowledge graph. It can be directly applied to various datasets.
@@
  • Adaptive-RAG [23] proposes an adaptive multi-step retrieval method that can dynamically select the most suitable retrieval strategy based on the complexity of the query.
  • FLARE [24] introduces a multi-step retrieval method that actively decide when and how to retrieve documents. It also predicts the future content to the guide the retrieval in next steps.
  • IRCoT [64] is a powerful multi-step retrieval pipeline that integrates the retrieval with the chain-of-thought (CoT) reasoning of LLMs. It guides the retrieval with CoT and in turn using retrieved documents to improve CoT. IRCoT can be compatible with arbitrary retrievers to conduct multi-step retrieval and reasoning.
Appendix D Implementations and Training Details
D.1 Training Data Construction
We extract 60,000 samples from the training set of HotpotQA, MuSiQu, and 2Wiki to construct KG-indexes and conduct large-scale training. Specifically, we merge the candidate passages as the document corpus. In the KG-index construction, we use the GPT-4o-mini [47] with the OpenIE prompts described in HippoRAG [16] to extract the entities, relations, and triples from the document corpus. Then, we use the ColBERTv2 [55] to conduct the entity resolution by computing the similarity between entities as
@@ D.2 Model Settings
Supervised Retrieval Fine-tuning $\alpha$ 0.3
@@
D.3 Training Settings
In KG completion pre-training, we randomly sample triples $(e,r,t)$ from knowledge graphs and mask out either the head or the tail entity to create a synthetic query $q=(e,r,?)$ and answer $a=\{e\}$ in a self-supervised manner. For example, given a triple (Barack Obama, born_in, Honolulu), we can create a query as $(\texttt{Barack Obama},\texttt{born\_in},?)$, which is encoded as a sentence embedding and fed into the GFM to predict the target entity Honolulu on graphs.
In supervised document retrieval fine-tuning, we obtain natural language questions and supporting documents from the multi-hop QA datasets. For each question, we identify the entities from its supporting documents as the targets. For instance, given the question “Where was Barack Obama born in?”, we can extract two entities such as [Honolulu, USA] from its supporting documents (e.g., Doc. 2 in Figure˜2). The GFM is trained to maximize the likelihood of these two target entities.
In the self-supervised KG completion pre-training, the GFM is trained on the mixture of 60 constructed KG-indexes for 30,000 steps. Then, we conduct the supervised document retrieval fine-tuning on the labeled question-document pairs for 5 epochs. The weight $\alpha$ between losses is set to 0.3. We use AdamW optimizer, learning rate of 5e-4 with batch sizes of both training stages set to 4. Each batch contains only one KG-index and training samples associated to it, where we randomly sample from different KG-indexes during training. The model is trained on 8 NVIDIA A100s (80G) with 14 hours pre-training and 5 hours supervised fine-tuning. The detailed settings are summarized in Table˜7.
Appendix E Additional Experiments
@@
Table 9: Comparison of GFM-RAG with pre-trained and fine-tuned sentence embedding models.
Method HotpotQA MuSiQue 2Wiki
R@2 R@5 R@2 R@5 R@2 R@5
GFM-RAG 78.3 87.1 49.1 58.2 90.8 95.6
all-mpnet-v2 (pre-trained) 59.4 73.3 33.2 46.3 48.5 59.4
all-mpnet-v2 (finetuned) 67.0 82.3 41.7 55.0 65.1 76.7
E.1 Effectiveness of Different Sentence Embeddings
In this section, we first study the effectiveness of different sentence embeddings in the GFM. We compare the all-mpnet-v2 [57], bge-large-en [69], gte-Qwen2-1.5B-instruct and gte-Qwen2-7B-instruct [34] as well as NV-Embed-v2 [31]. We download the official pre-trained model from the Huggingface [^1] https://huggingface.co/. The details of the models are shown in Table˜8. From the results, we can observe that the performance variance between different sentence embeddings is relatively small, where the all-mpnet-v2 achieves the best performance with respect to 3 metrics. This indicates that GFM-RAG is not sensitive to the choice of sentence embedding models. In experiments, we use the all-mpnet-v2 as the default sentence embedding model due to its efficiency. However, it has relative smaller context-size (512) which limits the length of input text. We leave the exploration of larger context-size sentence embedding models (e.g., NV-Embed-v2 with 32k context) for future work.
Then, we expand our ablation study to compare GFM-RAG with variants without GNN and using solely the pre-trained all-mpnet-v2 embeddings and those fine-tuned on multi-hop QA data, respectively. The results are shown in Table˜9. We can observe that GNN plays a crucial role in retrieval. The sentence embedding model all-mpnet-v2 is pre-trained on large-scale text data and could potentially see the QA data. However, it is not specifically trained for the multi-hop QA task, which leads to suboptimal performance in capturing the relationship between question and supporting documents. The fine-tuned all-mpnet-v2 achieves better performance than the pre-trained one by supervised fine-tuning on the multi-hop QA data, but still inferior to GFM-RAG. This indicates that the GNN can effectively capture the relationship between knowledge and conduct multi-hop reasoning, which is not achievable by simply using the sentence embedding model.
E.2 Effectiveness of Different Training Strategies
In this section, we first study the effectiveness of the two training tasks used in GFM-RAG. We compare the performance by only conducting the KG completion pre-training (GFM-RAG $w/o$ Fine-tune) and supervised document retrieval fine-tuning (GFM-RAG $w/o$ Pre-train). The results are shown in Table˜10. The results show that removing the supervised document retrieval fine-tuning significantly decreases the performance of GFM-RAG. This highlights the importance of supervised fine-tuning, as it enables the model to understand users’ queries and better capture the relevance between questions and knowledge for improved retrieval.
Table 10: Effectiveness of KGC pre-training and supervised retrieval fine-tuning in GFM-RAG.
Method HotpotQA MuSique 2Wiki
@@ E.2 Effectiveness of Different Training Strategies
GFM-RAG $w/o$ Retrieval Fine-tune 21.0 32.8 18.3 25.9 44.6 53.4
GFM-RAG $w/o$ KGC Pre-train 77.8 86.5 48.3 58.3 88.3 92.5
Table 11: Knowledge graph completion result of different training strategies.
Method MRR Hits@1 Hits@3 Hits@10
GFM-RAG 0.193 0.138 0.221 0.293
GFM-RAG $w/o$ Retrieval Fine-tune 0.304 0.234 0.323 0.451
GFM-RAG $w/o$ KGC Pre-train
0.029 0.007 0.022 0.067
Although the pre-training has a relatively small impact on the final performance, its primary purpose is to learn the general graph reasoning ability, following previous studies like ULTRA [11]. This would enhance the generalization and robustness of the GFM, which could be beneficial to its performance on other tasks, such as knowledge graph completion. To further validate this, we conduct an ablation study to compare GFM-RAG with different training strategies on the knowledge graph completion task. We report the knowledge graph completion (KGC) performance on the KG-index from the test set of the HotpotQA dataset. The results are shown in Table˜11.
From the knowledge graph completion results, we can observe that the GFM-RAG undergoes only the pre-training (GFM-RAG $w/o$ Fine-tune) achieves the best performance, which indicates that the pre-training is effective in learning the general graph reasoning ability. The performance of GFM-RAG with only supervised fine-tuning (GFM-RAG $w/o$ Pre-train) is significantly lower than that of GFM-RAG with pre-training. This indicates that the supervised fine-tuning is only learning the specific downstream task, which would limit the generalization ability of GFM-RAG as the foundation model. The GFM trained with both pre-training and supervised fine-tuning achieves the second-best performance on the knowledge graph completion task and the best performance on the multi-hop QA task. This indicates that both training strategies are essential for GFM-RAG to learn the general graph reasoning ability and benefit specific downstream tasks.
E.3 Effectiveness of Loss Weights
In this section, we examine the effectiveness of the weights assigned to the BCE loss and ranking loss in training GFM-RAG. We compare performance by varying the weight $\alpha$ between the two losses: ${\mathcal{L}}=\alpha{\mathcal{L}}_{\text{BCE}}+(1-\alpha){\mathcal{L}}_{\text{RANK}}$, with results presented in Table˜12. The findings indicate that using only either the BCE loss or ranking loss leads to suboptimal performance ($\alpha=0~\text{or}~1$). The best performance occurs when $\alpha$ is set to 0.3, which aligns with previous studies [36] suggesting that a smaller weight for BCE loss is preferable when positive samples are rare in the training data.
@@ E.3 Effectiveness of Loss Weights
E.4 Effectiveness of Ranking Methods
In this section, we investigate the effectiveness of different ranking methods based on inverted index used in GFM-RAG. We compare four ranking methods including (1) IDF + Top-T Pred: Our proposed method (eqs.˜14, 15 and 16), which maps the top-T entities predicted by GFM to documents using inverse document frequency (IDF)-weighted scores. (2) IDF + All Pred: Uses all predicted entities from GFM and weights them by IDF ($w/o$ eq.˜14). (3) Top-T Pred: Uses only the top-T predicted entities without applying IDF weighting ($w/o$ eq.˜15). (4) All Pred: Use all entity predictions and directly map to document scores ($w/o$ eqs.˜14 and 15). The results are shown in Table˜13. The results show that the proposed IDF + Top-k Pred performs the best. This indicates that the inverted index is a crucial component of GFM-RAG, which serves as a bridge between structured reasoning over KGs and the unstructured documents required by LLMs, necessitating a careful design.
We acknowledge the potential alternatives, and as a promising future direction, we plan to explore end-to-end models that can jointly reason over structured and unstructured knowledge without relying on an explicit inverted index.
Table 13: Comparison of different ranking methods.
Ranking Method HotpotQA MuSiQue 2Wiki
R@2 R@5 R@2 R@5 R@2 R@5
IDF + Top-T Pred (GFM-RAG) 78.3 87.1 49.1 58.2 90.8 95.6
IDF + All Pred ($w/o$ eq.˜14) 68.1 71.4 35.8 41.2 86.0 87.5
Top-T Pred ($w/o$ eq.˜15) 71.6 78.6 46.3 52.5 74.7 78.1
All Pred ($w/o$ eqs.˜14 and 15) 77.6 82.9 41.1 46.9 88.6 90.4
E.5 Ablation Study of Training Datasets
We further conducted ablation studies where GFM-RAG is trained separately on each dataset, and we report performance across all three benchmarks. Results are shown Table˜14. These results show that GFM-RAG not only performs well on the trained datasets, but also generalizes well to other datasets. More importantly, the model trained on multi-domain datasets performs competitively across all datasets, validating its ability to generalize effectively across domains and benefit from training on diverse KGs by learning generalizable reasoning ability across domains.
Table 14: Ablation study of GFM-RAG trained on each dataset. Best results are highlighted in bold. The second best is underlined.
Test Dataset HotpotQA MuSiQue 2Wiki
Training Dataset R@2 R@5 R@2 R@5 R@2 R@5
HotpotQA 79.3 87.8 46.9 57.2 86.6 92.4
MusiQue 68.8 81.8 47.6 57.5 84.4 89.6
2Wiki 72.2 77.9 46.6 55.5 89.3 93.2
All 78.3 87.1 49.1 58.2 90.8 95.6
E.6 Model Transferability
@@
GFM-RAG (domain-specific fine-tuning) 82.7 75.9 60.8 49.5 77.5 86.6
In this section, we evaluate GFM-RAG’s transferability by conducting domain-specific fine-tuning on the training split of dataset on each domain. As shown in 15, GFM-RAG performs well in zero-shot generalization, with further improvements achieved through domain-specific fine-tuning. This highlights its transferability when adapted to domain-specific datasets.
@@
E.7 Details of Model Neural Scaling
In this section, we provide more details on the neural scaling experiments. We evaluate the changes of the model performance with respect to different parameter sizes and training data sizes. In GFM-RAG, the model parameter sizes are primarily influenced by the hidden dimension of the GFM. Thus, we vary the dimension from 32 to 512 which results in the model parameter sizes ranging from 0.08M to 8M. The detailed settings are shown in Table˜16. We test models with different sizes on different scales of training data ranging from 3k to 45k samples. We separately report the fitted trend line of performance changing with model parameter size and training data size in Figure˜5. From the trend line, we can observe that the performance of GFM-RAG increases with the model parameter size and training data size. Meanwhile, with the larger model parameter size a larger training data size is required to achieve the best performance. This indicates that the performance of GFM-RAG can be further improved by scaling up the model size and training data simultaneously.
To further investigate architectural design, we varied the number of GNN layers from 1 to 8 while keeping the hidden dimension fixed (512), and evaluated model performance across all datasets. The results are shown in Table˜17. We observe that performance generally improves with deeper GNN layers, which we attribute to both the increased model sizes and the ability to capture more complex multi-hop associations. This trend aligns with the neural scaling laws observed in foundation models, where larger parameter counts typically yield better generalization.
Interestingly, we find that performance peaks around 4 layers in some cases. As discussed in Appendix˜A and Section˜4.8, GFM-RAG is designed to capture logical associations from KGs through multi-hop message passing. However, since the maximum number of reasoning hops required by our datasets is 4, additional layers beyond this offer limited benefit, likely due to the absence of higher-hop training signals. This finding supports our hypothesis that GFM-RAG effectively learns query-relevant multi-hop reasoning paths, and that deeper architectures may not improve performance without datasets requiring more complex reasoning. In summary, these results demonstrate the effectiveness and interpretability of the proposed GNN-based architecture, and confirm that both model capacity and logical expressibility contribute to GFM-RAG ’s strong performance. We recognize the potential of other architectural designs and aim to explore them in the future, inspiring the community to do the same.